the cold and hot water faucets of a certain bathtub take 7 minutes to fill the tub. If it takes the hot water faucet 17 minutes to fill the tub by itself, how long will it take the cold water faucet to fill the tub on its own?
Answer by sachi(548)

let the cold water faucet takes x min to fill the tub on its own
so in 1 min it will fill 1/x part
the hot water faucet takes 17 minutes to fill the tub by itself
so in 1 min it will fill 1/17 part
so in 1 min the cold and hot water faucets will fill (1/x+1/17) part
but the cold and hot water faucets take 7 minutes to fill the tub
so 1/x+1/17=1/7
or 1/x=1/7-1/17=(17-7)/17*7=10/119
or x=119/10=11.9 min
so the cold water faucet takes 11.9 min to fill the tub on its own
ans